OpenStudy (anonymous):

Simplify. Write in scientific notation. (9 × 10^–3)^2

OpenStudy (anonymous):

12345.67901 or in scientific, it is equal to: 1.234567901*10^5

OpenStudy (anonymous):

were did all those numbers come from??

OpenStudy (anonymous):

is it 8.1 x 10^-5

OpenStudy (anonymous):

Sorry, i read that as a -2 in the end, and yea, what you got is right.

OpenStudy (anonymous):

okay good :) now here is the most hard one.\[Evaluate \frac{ 1 }{ 2} a ^{-4} b ^{2} for a = –2 and b = 4.\]

OpenStudy (anonymous):

What did I say about using the equation editor? Math ONLY. Anyhow, just plug in for the calculator and you get: \[-.2\] or \[-\frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

neither of those are my options. 2 \[\frac{ 1 }{ 2 }\] 3 4

OpenStudy (anonymous):

Hmm, can I see the original question then? Like a pic?

OpenStudy (anonymous):

Actually, I see. Your writing messed me up again. It should be 1/2.

OpenStudy (anonymous):

8. Is the following statement always, never, or sometimes true? A number raised to a negative exponent is negative. always never sometimes